# Conquering the Communication Constraints to Enable Large Pre-Trained Models in Federated Learning

### v3

@@ Abstract
Federated learning (FL) has emerged as a promising paradigm for enabling the collaborative training of models without centralized access to the raw data on local devices. In the typical FL paradigm (e.g., FedAvg), model weights are sent to and from the server each round to participating clients. Recently, the use of small pre-trained models has been shown to be effective in federated learning optimization and improving convergence. However, recent state-of-the-art pre-trained models are getting more capable but also have more parameters. In conventional FL, sharing the enormous model weights can quickly put a massive communication burden on the system, especially if more capable models are employed. Can we find a solution to enable those strong and readily available pre-trained models in FL to achieve excellent performance while simultaneously reducing the communication burden? To this end, we investigate the use of parameter-efficient fine-tuning in federated learning and thus introduce a new framework: FedPEFT. Specifically, we systemically evaluate the performance of FedPEFT across a variety of client stability, data distribution, and differential privacy settings. By only locally tuning and globally sharing a small portion of the model weights, significant reductions in the total communication overhead can be achieved while maintaining competitive or even better performance in a wide range of federated learning scenarios, providing insight into a new paradigm for practical and effective federated systems. Code is available at https://github.com/imguangyu/FedPEFT.
@@
Federated learning, Parameter-efficient fine-tuning, Vision transformers, Image Classification, Action Recognition.
Figure 1: Process in a federated learning communication round with $\mathbf{M}$ participating clients. We use ViT-Base as an instance to analyze the communication costs. (a) Conventional federated learning framework, where the entire model will be sent during the communication. (b) FedPEFT, which is our proposed parameter-efficient framework for federated learning. Instead of uploading and downloading the entire model during the communication, only a small subset of the model will be updated and communicated, leading to a huge relief in communication, specifically from about 328MB per client per round to about 0.68MB per client per round.
@@ I Introduction
When employing strong pre-trained models, the number of parameters can be large, such as for current state-of-the-art transformers. For example, ViT-Base (ViT-B) [11] has 84 million parameters, let alone the current significant progress in large foundation models (e.g., GPT-4 [12] has 1 trillion parameters). Those large models would simply exacerbate the communication overhead to insurmountable levels. As a compromise, most existing FL work focuses on the performance of smaller Convolutional Neural Networks (e.g., ResNet [13]) on smaller datasets (e.g., CIFAR-10 [14], EMINIST [15]). Considering the thriving progress in large pre-trained Foundation Models [16], an efficient framework enabling these large pre-trained models will be significant for the FL community.
@@
Can we reap these potential communication benefits while still achieving strong performance in FL? Unfortunately, operating conditions in FL are difficult, requiring successful convergence under varying data heterogeneity levels, random client availability, and differential privacy procedures. Therefore, we are unable to properly assess this possibility of benefit based on existing literature, as the effectiveness of parameter-efficient fine-tuning methods has not been explored in such situations in FL. To fill this gap, we explore the viability of a Federated Parameter-Efficient Fine-Tuning (FedPEFT) framework with a systemic empirical study on a comprehensive set of FL scenarios including communication analysis about communication cost for each method to enable pre-trained models, capability analysis of each method with unlimited communication budget, and robustness analysis of each method when additional constraints (i.e., differential privacy or data scarcity) applied. The framework is illustrated in Fig. 1. We deploy parameter-efficient fine-tuning methods to adapt pre-trained models and enable massive reductions in communication overheads.
@@
  • We introduce FedPEFT, a new federated learning framework that simultaneously addresses data heterogeneity and communication challenges. FedPEFT is the first federated learning framework that enables the leveraging of strong pre-trained models in FL while maintaining an extremely low communication cost.
@@
  • To ensure FedPEFT is practical for the complex environments of FL, we further analyze the robustness of FedPEFT among low-data regimes, differential privacy operations, and an ablation study with different pre-trained backbones. (Sections IV-D and V-A)
@@ II Related Work
Parameter-Efficient Fine-tuning. Fine-tuning is a prevalent topic in centralized transfer learning, especially in this era of the “Foundation Model” [16]. A significant line of work is to reduce the trainable parameter number, i.e., parameter-efficient fine-tuning (PEFT) [20, 27, 28, 29, 30, 31, 17, 32]. PEFT has emerged as a pivotal area of research in the field of natural language processing (NLP) [19, 33, 34, 35] and further adapted into more fields such as computer vision (CV) [18, 36, 37, 38, 39, 40, 41]. With different focuses, PEFT methods can be divided into three categories: 1) Input Adaptation methods such as prompt-tuning [18, 30] focusing on adding learnable context to the input data, 2) Backbone Adaptation methods such as adapter-tuning [38, 19, 33], and 3) Specification methods such as bias-tuning [17, 31]. PEFT enables easier access and usage of pre-trained models by reducing the memory cost needed to conduct fine-tuning due to fewer computed gradients. In federated learning, PEFT has an additional benefit that is not salient in centralized training: reducing communication costs. By introducing PEFT to federated learning, our work can take advantage of a strong (and even large) pre-trained model while significantly reducing communication costs. Several concurrent and follow-up works study the PEFT methods under FL settings in NLP [42, 43, 44, 45]. Complementarily, our framework is the first work to present a comprehensive study of PEFT in various FL settings in computer vision under both image and video tasks.
@@ III-A Problem Formulation
Input: $N$ clients indexed by $i$, participating-client number $M$, communication rounds $T$, trainable parameters $\delta$ of the model where $|\delta|<<|\phi|$, pre-trained model weights ${\bm{\theta}}$, random initialized $\delta^{(0)}$, and local epoch number $E$. Server executes:       initialize $\phi^{(0)}$ and $\{\phi_{i}^{(0)}\}_{i=1}^{N}$ with ${\bm{\theta}}$ and $\delta^{(0)}$       for each round $t=0,2,...,T-1$ do        ${\mathbb{S}}_{t}\leftarrow$ (random set of $M$ clients)        for each client $i\in{\mathbb{S}}_{t}$ in parallel do         $\delta_{i}^{(t)}\leftarrow\delta^{(t)}$         $\delta_{i}^{(t+1)}\leftarrow\textbf{ClientUpdate}(\delta_{i}^{(t)},i)$        $\delta^{(t+1)}=\sum_{i=1}^{M}\frac{|{\mathbb{D}}_{i}|}{\sum_{i=1}^{M}|{\mathbb{D}}_{i}|}\delta_{i}^{(t+1)}$        $\phi^{(t+1)}\leftarrow\{{\bm{\theta}},\delta^{(t+1)}\}$       return $\phi^{(T)}$ ClientUpdate ($\delta,i$):       $\delta\leftarrow$ perform local training on $\delta$ with ${\mathbb{D}}_{i}$ for $E$ epochs       return $\delta$
Figure 2: Methods to fine-tune each layer in a pre-trained backbone, where $h$ means the input, $\phi$ means the pre-trained layer, and $\phi_{w},\phi_{b}$ mean its weight and bias parameters, respectively. (a) Fully fine-tune the entire model. (b) Only tune the classification head with the frozen backbone. Blue and red rectangles mean frozen and trainable parameters, respectively. (c) Only tune the bias term of each layer. (d) Concatenate learnable parameters $\phi_{p}$ as prompts to the input in each layer and only tune the prompts with the frozen backbone. (e) Add an adapter $\phi_{a}$ to each layer of the backbone, and only tune the adapters with the frozen backbone.
III-B FedPEFT
@@
To implement FedPEFT, we provide a canonical baseline approach (Head-tuning) and three prototypes leveraging different parameter-efficient fine-tuning methods (Bias, Adapter, and Prompt), which are detailed in the following.
@@
FedPEFT-Bias. Bias-tuning [17] aims to adapt the pre-trained backbone by only fine-tuning a specific group of parameters, the bias term. In this way, the backbone can be trained with moderate adjustments to prevent damaging the upstream representation. We show the output of each layer $\phi^{l}$ given hidden states $h$ as
[…]
where $\phi_{w}^{L}$ and $\phi_{b}^{L}$ are weight and bias of $\phi_{L}$, blue parameters are frozen, while red parameters are trainable.
FedPEFT-Adapter. Instead of directly tuning existing parameters in the backbone like Bias-tuning, Adapter-tuning [19] adds a few parameters called adapters inside the backbone $\phi$ instead. Usually, adapters will be deployed in each layer of the backbone to perform transformations on different levels of the pre-trained feature while the backbone stays frozen. We show the output of each layer $\phi^{l}$ given hidden states $h$ as
[…]
where $\phi_{a}$ is the adapter.
FedPEFT-Prompt. Prompt-tuning [18] takes a slightly different approach from the other fine-tuning methods. Specifically, it concatenates trainable parameters, called prompt embeddings, to the input embedding and hidden states in each layer. We show the output of each layer $\phi^{l}$ given hidden states $h$ as
@@
We illustrate the differences between all baseline and prototype methods in Fig. 2. Besides the above prototypes, our framework is compatible with other PEFT methods. We provide additional discussion on Sec. V-B.
@@ III-D Privacy Discussion
TABLE I: Communication analysis on the image domain. The communication cost is computed with 4B/parameter. The averaged final accuracy, i.e., $t=T=50$, and the standard deviation of three different random seeds are reported for each data set. The number of tuned parameters is computed based on CIFAR-100, but it may be slightly different for each dataset. The first section shows the change in accuracy when decreasing the participating client number. The gray numbers indicate the baseline performance with no decrease in the participating client number. The second section shows the change in accuracy when we reward the low communication cost of head-tuning by increasing the number of participating clients. The third section shows the accuracy when we fully fine-tune a lightweight model, ShuffleNet V2 0.5×0.5$\times$0.5 × [50]. The fourth section shows the performance of each prototype of FedPEFT.
@@ IV Experiments
Communication Analysis: When faced with a limited communication budget, there are several conventional solutions to reduce costs, e.g., sampling fewer clients each round or using a lightweight model. Can FedPEFT outperform other solutions in terms of communication cost and accuracy? (RQ1)
@@ IV-A Experiments details
Dataset. For our study, we focus on computer vision (CV) applications as our testbed. Specifically, we investigate the performance of each method on the Image and Video domains with image classification and action recognization tasks. For the image classification, we employ ImageNet-21K [51] as the pre-training dataset, which is a primary dataset for pre-training models in CV. Then we select three datasets for the downstream tasks that have increasing degrees of domain gap compared to ImageNet-21k, and we visualize and quantify the domain gap in Section IV-C: Resisc45 [52], CIFAR-100 [14] and PCam [53]. Resisc45 is a remote sensing image dataset containing 18,900 training samples and 6,300 testing samples for 45 scene classes, which has a small domain gap with ImangeNet-21K. CIFAR-100 contains 50,000 training samples and 10,000 testing samples for 100 classes of nature objects (e.g., sunflowers, buses), which has a medium domain gap. PCam is a medical dataset containing 32,768 colored images extracted from histopathologic scans of lymph node sections, which has a much larger domain gap. Each image is annotated with a binary label indicating the presence of metastatic tissue. We sample 20,000 images from the original train set for training and use the entire test set for evaluation. We use the official splitting for CIFAR-100 and PCam and a public splitting from AiTLAS [54] for Resisc45.
For action recognition evaluation, we employ Kinetics-400 [55] as the pre-training dataset, which is a common dataset for pretraining in action recognition. We select three datasets with varying degrees of domain gap as compared to Kinetics-400: UCF101 [56], HMDB51 [57] and UCF-CRIME [58]. The UCF101 [56] dataset contains 13,320 videos from 101 classes. The HMDB51 [57] dataset contains 6849 videos divided into 51 categories based on everyday human actions. UCF-CRIME [58] consists of surveillance videos with 13 anomalous activities as action recognition classes. For UCF-CRIME, we use a train/val split of 40/10 for each class.
@@
Implementation Detail. We choose ViT-B [11] with image size 224 and patch size 16 as our backbone. For the image domain, the backbone is pre-trained on ImageNet-21K [51], as available in the timm library [59]. The images for the downstream datasets are resized to $224\times 224$. Images from CIFAR-100 are augmented by random cropping with a padding of $4$ and random horizontal flipping, and Resisc45 and PCam are augmented only with random horizontal flipping. For the video domain, we choose vanilla ViT-Base (ViT-B/16) [11] with joint space-time attention as our backbone model [60] using VideoMAE [61] pre-trained weights on Kinetics-400 [55]. We perform the experiments on 8 Nvidia RTX A5000 GPUs with a batch size of $64$. All reported numbers are run multiple times and averaged. The base hyperparameters for each method are described below.
  • Full fine-tuning: We use SGD [62] optimizer with learning rate $0.001$ and weight decay $0.0001$.
  • Head-tuning: We use one linear layer as the classification head and use the SGD optimizer with learning rate $0.005$ and weight decay $0.0001$.
  • FedPEFT-Bias: There is no additional hyperparameter for Bias. We use SGD optimizer with learning rate $0.01$ and weight decay $0.0001$.
  • FedPEFT-Adapter: We use a Bottleneck Adapter with residual connections as the adapter. We insert the adapter to each layer of the backbone after the feed-forward block following [19] with a reduction factor of $8$ and the GELU as the activation function. We use SGD optimizer with learning rate $0.005$ and weight decay $0.0001$.
  • FedPEFT-Prompt: We follow the design of VPT-Deep [18], and use a prompt length as $10$ to prepend prompt tokens to the input and the hidden embedding in each layer. We use SGD optimizer with learning rate $0.01$ and weight decay $0.0001$.
@@ IV-B RQ1: Communication Analysis
To verify the effectiveness of FedPEFT and answer the first research question (RQ1, Section IV), we compare it with three baselines while monitoring the communication budget: a) Full fine-tuning of our default model (ViT-B). We vary the number of participating clients to show different levels of communication requirements. b) Head-tuning. The communication cost of head-tuning is naturally lower than other methods, so we increase the participating clients to make it a stronger baseline. c) Fully fine-tune a light-weighted model (ShuffleNet V2 0.5$\times$ [50] for images and X3D-S [63] for videos) with a similar parameter scale.
As demonstrated in TABLE I for the image domain and TABLE II for the video domain, all FedPEFT methods achieve better results in many cases compared with other approaches, even with significantly fewer communicated parameters. We find that full fine-tuning needs several orders of magnitude of communication to achieve a comparable result with FedPEFT. For instance, it needs at least 187$\times$ and 477$\times$ more parameters to reach and outperform FedPEFT on CIFAR-100. Interestingly, full fine-tuning performs well on Resisc45 where the domain gap is smaller, even when the participating-client number is low. However, when the domain gap increases, more participating clients will be needed to outperform FedPEFT, and finally it fails to outperform FedPEFT even without reducing the participating-client number on PCam where a large domain gap exists. Meanwhile, head-tuning lags behind most other approaches, but the performance is stable with different levels of domain gap, while the ShuffleNet model only achieves $71\%$, $55\%$, and $89\%$ of accuracy on Resisc45, CIFAR-100, and PCam with 2.4$\times$ the communication cost compared with FedPEFT-Bias. Besides, the standard deviation of FedPEFT is lower than most other solutions, especially when the domain gap is large, showing the stability of FedPEFT. For the video domain, the conclusion is consistent.
@@
Figure 4: Visualization and analysis of domain gap. (a) We apply Linear Discriminant Analysis (LDA) to the extracted feature from the pre-trained backbone to visualize and quantify the domain gap. (b) We plot the performance (relative accuracy) of each PEFT method compared against full fine-tuning. The gray dashed line indicates the performance of full fine-tuning (i.e., reference), and the three points in each line are the relative server accuracy on Resisc45, CIFAR-100, and PCam, respectively.
TABLE III: Capability analysis among different federated learning settings on CIFAR-100 and UCF-101. Training from scratch in more complicated settings will lead to a lower result than in the first setting, which is omitted here. Bold-style shows the best performance among all methods or among prototypes in FedFEFT.

### v5

# Exploring Parameter-Efficient Fine-Tuning to Enable Foundation Models in Federated Learning

@@ Abstract
Federated learning (FL) has emerged as a promising paradigm for enabling the collaborative training of models without centralized access to the raw data on local devices. In the typical FL paradigm (e.g., FedAvg), model weights are sent to and from the server each round to participating clients. Recently, the use of small pre-trained models has been shown to be effective in federated learning optimization and improving convergence. However, recent state-of-the-art pre-trained models are getting more capable but also have more parameters, known as the “Foundation Models.” In conventional FL, sharing the enormous model weights can quickly put a massive communication burden on the system, especially if more capable models are employed. Can we find a solution to enable those strong and readily available pre-trained models in FL to achieve excellent performance while simultaneously reducing the communication burden? To this end, we investigate the use of parameter-efficient fine-tuning in federated learning and thus introduce a new framework: FedPEFT. Specifically, we systemically evaluate the performance of FedPEFT across a variety of client stability, data distribution, and differential privacy settings. By only locally tuning and globally sharing a small portion of the model weights, significant reductions in the total communication overhead can be achieved while maintaining competitive or even better performance in a wide range of federated learning scenarios, providing insight into a new paradigm for practical and effective federated systems.
@@
federated learning, parameter-efficient fine-tuning, vision transformers, image classification, action recognition
@@ I Introduction
When employing strong pre-trained models, the number of parameters can be large, such as for current state-of-the-art transformers. For example, ViT-Base (ViT-B) [11] has 84 million parameters, let alone the current significant progress in large foundation models (e.g., GPT-4 [12] has more than 1 trillion parameters). Those large models would simply exacerbate the communication overhead to insurmountable levels. As a compromise, most existing FL work focuses on the performance of smaller Convolutional Neural Networks (e.g., ResNet [13]) on smaller datasets (e.g., CIFAR-10 [14], EMINIST [15]). Considering the thriving progress in large pre-trained Foundation Models [16], an efficient framework enabling these large pre-trained models will be significant for the FL community.
Figure 1: Process in a federated learning communication round with $\mathbf{M}$ participating clients. We use ViT-Base as an instance to analyze the communication costs. (a) Conventional federated learning framework, where the entire model will be sent during the communication. (b) FedPEFT, which is our proposed parameter-efficient framework for federated learning.
@@
Can we reap these potential communication benefits while still achieving strong performance in FL? Unfortunately, operating conditions in FL are difficult, requiring successful convergence under varying data heterogeneity levels, random client availability, and differential privacy procedures. Therefore, we are unable to properly assess this possibility of benefit based on existing literature, as diverse parameter-efficient fine-tuning methods have not been systematically explored in such situations in FL. To fill this gap, we explore the viability of a Federated Parameter-Efficient Fine-Tuning (FedPEFT) framework with a systemic empirical study on a comprehensive set of FL scenarios including communication analysis about communication cost for each method to enable pre-trained models, capability analysis of each method with unlimited communication budget, and robustness analysis of each method when additional constraints (i.e., differential privacy or data scarcity) applied. The framework is illustrated in Fig. 1. We deploy parameter-efficient fine-tuning methods to adapt pre-trained models and enable massive reductions in communication overheads.
@@
  • We explored several PEFT methods in FL as the FedPEFT framework to simultaneously addresses data heterogeneity and communication challenges. FedPEFT allows for the utilization of powerful pre-trained models in federated learning while keeping communication costs extremely low.
@@
  • To ensure FedPEFT is practical for the complex environments of FL, we further analyze the robustness of FedPEFT among low-data regimes and differential privacy operations. (Sections IV-D)
@@ II Related Work
Parameter-Efficient Fine-tuning. Fine-tuning is a prevalent topic in centralized transfer learning, especially in this era of the “Foundation Model” [16]. A significant line of work is to reduce the trainable parameter number, i.e., parameter-efficient fine-tuning (PEFT) [20, 27, 28, 29, 30, 31, 17, 32]. PEFT has emerged as a pivotal area of research in the field of natural language processing (NLP) [19, 33, 34, 35] and further adapted into more fields such as computer vision (CV) [18, 36, 37, 38, 39, 40, 41]. With different focuses, PEFT methods can be divided into three categories: 1) Input Adaptation methods such as prompt-tuning [18, 30] focusing on adding learnable context to the input data, 2) Backbone Adaptation methods such as adapter-tuning [38, 19, 33], and 3) Specification methods such as bias-tuning [17, 31]. PEFT enables easier access and usage of pre-trained models by reducing the memory cost needed to conduct fine-tuning due to fewer computed gradients. In federated learning, PEFT has an additional benefit that is not salient in centralized training: reducing communication costs. By introducing PEFT to federated learning, our work can take advantage of a strong (and even large) pre-trained model while significantly reducing communication costs. Several works study the PEFT methods under FL settings in NLP [42, 43, 44, 45]. Complementarily, our work provides a comprehensive study of PEFT in various FL settings in computer vision under both image and video tasks and insights on various scenarios, including more privacy requirements or under limited data.
@@ III-A Problem Formulation
III-B FedPEFT
Figure 2: Methods to fine-tune each layer in a pre-trained backbone, where $h$ means the input, $\phi$ means the pre-trained layer, and $\phi_{w},\phi_{b}$ mean its weight and bias parameters, respectively.
Input: $N$ clients indexed by $i$, participating-client number $M$, communication rounds $T$, trainable parameters $\delta$ of the model where $|\delta|<<|\phi|$, pre-trained model weights ${\bm{\theta}}$, random initialized $\delta^{(0)}$, and local epoch number $E$.
Server executes:       initialize $\phi^{(0)}$ and $\{\phi_{i}^{(0)}\}_{i=1}^{N}$ with ${\bm{\theta}}$ and $\delta^{(0)}$       for each round $t=0,2,...,T-1$ do        ${\mathbb{S}}_{t}\leftarrow$ (random set of $M$ clients)        for each client $i\in{\mathbb{S}}_{t}$ in parallel do         $\delta_{i}^{(t)}\leftarrow\delta^{(t)}$         $\delta_{i}^{(t+1)}\leftarrow\textbf{ClientUpdate}(\delta_{i}^{(t)},i)$        $\delta^{(t+1)}=\sum_{i=1}^{M}\frac{|{\mathbb{D}}_{i}|}{\sum_{i=1}^{M}|{\mathbb{D}}_{i}|}\delta_{i}^{(t+1)}$        $\phi^{(t+1)}\leftarrow\{{\bm{\theta}},\delta^{(t+1)}\}$       return $\phi^{(T)}$ ClientUpdate ($\delta,i$):       $\delta\leftarrow$ perform local training on $\delta$ with ${\mathbb{D}}_{i}$ for $E$ epochs       return $\delta$
@@ III-B FedPEFT
To implement FedPEFT, we provide a canonical baseline approach (Head-tuning) and three prototypes leveraging different representative parameter-efficient fine-tuning methods (Bias, Adapter, and Prompt), which are detailed in the following.
@@
[…]
We illustrate the differences between all baseline and prototype methods in Fig. 2. Besides the above prototypes, our framework is compatible with other PEFT methods such as LoRA [29].
@@ III-D Privacy Discussion
TABLE I: Communication analysis on the image (upper) and video (lower) domains. The communication cost is computed with 4B/parameter. The averaged final accuracy, i.e., $t=T=50$, and the standard deviation of three different random seeds are reported for each data set. The number of tuned parameters is computed based on CIFAR-100, but it may be slightly different for each dataset. The first section shows the change in accuracy when decreasing the participating client number. The gray numbers indicate the baseline performance with no decrease in the participating client number. The second section shows the change in accuracy when we reward the low communication cost of head-tuning by increasing the number of participating clients. The third section shows the accuracy when we fully fine-tune a lightweight model, ShuffleNet V2 0.5×0.5$\times$0.5 × [50] for image domain or X3D-S [51] for video domain. The fourth section shows the performance of each prototype of FedPEFT.
@@ IV Experiments
Communication Analysis: When faced with a limited communication budget, there are several solutions to reduce costs, e.g., sampling fewer clients each round or using a lightweight model. Can FedPEFT outperform other solutions in terms of communication cost and accuracy? (RQ1)
@@ IV-A Experiments details
Dataset. For our study, we focus on computer vision (CV) applications as our testbed. Specifically, we investigate the performance of each method on the Image and Video domains with image classification and action recognition tasks. For the image classification, we employ ImageNet-21K [52] as the pre-training dataset. Then we select three datasets for the downstream tasks that have increasing degrees of domain gap compared to ImageNet-21k, and we visualize and quantify the domain gap in Section 4: Resisc45 [53], CIFAR-100 [14] and PCam [54]. For video domain analysis, we take video action recognition task for evaluation. we employ Kinetics-400 [55] as the pre-training dataset and select three datasets with varying degrees of domain gap as compared to Kinetics-400: UCF101 [56], HMDB51 [57] and UCF-CRIME [58].
@@
Implementation Detail. We choose ViT-B [11] with image size 224 and patch size 16 as our backbone. For the image domain, the backbone is pre-trained on ImageNet-21K [52], as available in the timm library [59]. The images for the downstream datasets are resized to $224\times 224$. Images from CIFAR-100 are augmented by random cropping with a padding of $4$ and random horizontal flipping, and Resisc45 and PCam are augmented only with random horizontal flipping. For the video domain, we choose vanilla ViT-Base (ViT-B/16) [11] with joint space-time attention as our backbone model [60] using VideoMAE [61] pre-trained weights on Kinetics-400 [55]. We perform the experiments on 8 Nvidia RTX A5000 GPUs with a batch size of $64$. All reported main results are run under $3$ random seeds and averaged.
@@ IV-B RQ1: Communication Analysis
To verify the effectiveness of FedPEFT and answer the first research question (RQ1, Section IV), we compare it with three baselines while monitoring the communication budget: a) Full fine-tuning of our default model (ViT-B). We vary the number of participating clients to show different levels of communication requirements. b) Head-tuning. The communication cost of head-tuning is naturally lower than other methods, so we increase the participating clients to make it a stronger baseline. c) Fully fine-tune a light-weighted model (ShuffleNet V2 0.5$\times$ [50] for images and X3D-S [51] for videos) with a similar communication overhead.
As demonstrated in Table I, all FedPEFT methods achieve better results in many cases compared with other approaches, even with significantly fewer communicated parameters. We find that full fine-tuning needs several orders of magnitude of communication to achieve a comparable result with FedPEFT. For instance, it needs at least 187$\times$ and 477$\times$ more parameters to reach and outperform FedPEFT on CIFAR-100. Interestingly, full fine-tuning performs well on Resisc45 where the domain gap is smaller, even when the participating-client number is low. However, when the domain gap increases, more participating clients will be needed to outperform FedPEFT, and finally it fails to outperform FedPEFT even without reducing the participating-client number on PCam where a large domain gap exists. Meanwhile, head-tuning lags behind most other approaches, but the performance is stable with different levels of domain gap, while the ShuffleNet model only achieves $71\%$, $55\%$, and $89\%$ of accuracy on Resisc45, CIFAR-100, and PCam with 2.4$\times$ the communication cost compared with FedPEFT-Bias. Besides, the standard deviation of FedPEFT is lower than most other solutions, especially when the domain gap is large, showing the stability of FedPEFT. For the video domain, the conclusion is consistent.
@@
TABLE II: Capability analysis for different federated learning settings on CIFAR-100 and UCF-101. Training from scratch in more complicated settings will lead to a lower result than in the first setting, which is omitted here. Bold-style shows the best performance among all methods or among prototypes in FedFEFT.